40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  CTN Lab elog  Not logged in ELOG logo
Entry  Wed Feb 7 21:47:44 2018, Craig, awade, DailyProgress, NoiseBudget, February 7th Noisebudget 20180207_221703noiseBudget.pdf
    Reply  Thu Feb 8 18:38:04 2018, rana, DailyProgress, NoiseBudget, February 7th Noisebudget 
       Reply  Fri Feb 9 11:36:08 2018, awade, DailyProgress, NoiseBudget, 200-3000 Hz hump search 
          Reply  Mon Feb 12 16:29:50 2018, awade, Craig, DailyProgress, NoiseBudget, 200-3000 Hz hump search 
             Reply  Fri Feb 16 12:41:32 2018, awade, Craig, DailyProgress, NoiseBudget, 200-3000 Hz hump search Beatnote_ASD_20180215_121922.pdfBeatnote_ASD_20180215_193422.pdfBeatnote_CurrentASD.pdf
Message ID: 2083     Entry time: Fri Feb 9 11:36:08 2018     In reply to: 2082     Reply to this: 2087
Author: awade 
Type: DailyProgress 
Category: NoiseBudget 
Subject: 200-3000 Hz hump search 

So  1-2 kHz feature has popped up from time to time in previous beats.  It seems to be mostly arise from changes to the AEOM (installing) and poor polarization alignment of the RF EOMs and BB EOMS.  I can't find in the elog anywhere where it was explicitly pined down by tara optimized it out with fine tuning alignment of RF EOM (see PSL:1311).  Evan also made a measurement of the coherence between the RFAM (PSL:1524) and the beat, this looks like a pretty credible source of the hump.

My suspicion is that this noise hump is too high to be seismic (unless its unconverted).  RIN coupling in through photo thermal effects should dominate at lower frequencies unless the power spectrum of intensity noise is strongly skewed to  higher frequency components.  I've attached the Photo thermal transfer function from the noise budget notebook below for reference.  This has a trace with previous RIN measured in the old setup. We need  to measure RIN again, however, unless its very different its unlikely the source. 

Hypothesis #1:

The the RFAM out of the resonant modulation EOMs is poorly optimized.  The stationary component of this residual AM creates a DC offset in the error signal that isn't wholly nulled by the FSS offset adjust. This lock point error means that the actual lock point of the FSS is slightly down the cavity fringe and that any small fluctuations in intensity vary the slope of the error signal are directly coupled to first order as if they were frequency noise.  If the error signal is properly zeroed on the resonance we would be immune to this effect to first order.  Slow temperature variations may slightly change the polarizations so optimizing the voltage offset in the FSS may not hold true after a room temperature change.  In the past (PSL:1994) we had suppressed the AM component 80 dBm/sqrtHz when measuring on the reflection PD using a flipper mirror in front of the cavity.  It had been better that this in the past and is worse in the north path by an order of magnitude. This has not been checked in the last few months.

Hypothesis #2:

We are not just looking at a stationary lock point error from RFAM but also a component that originates from the FSS EOM path or is coupled back through it.  The BBEOM in the FSS loop can induce intensity changes if it not a optimal polarization alignment.  This may convert what is supposed to be a phase correction to the light into an intensity change.  If there is noise in the EOM path of the FSS this can convert into noise through hypothesis #1.  There could also be a feed back where the RIN is made worse by dynamically feeding back through the FSS making more RIN from the BBEOM.  

One possibility here to test this hypothesis is to inject a signal in the BBEOM path to see if we can induce a narrow peak (or possible band passed noise component) that couples all the way through to the BN and the FSS error spectrum. 

 

Not that the ISS is not currently turned on.  The optics need a quick tune up and the loops should be switched back on. The ISS was inducing some phase as part of the AEOMs operation.  You can see this in the actuation signal of the FSS as it fights to compensate the ISS.  Between the two of them they should suppress intensity noise and the higher frequency discrepancy with the reference cavity frequency: but, they are not entirely orthogonal in practice.  The ISS definitly make the FSS EOM work harder.

Some measurements to make

  • Free running RIN of both the north and south paths
  • RIN when ISS activated and comparison of BN spectrum change
  • RFAM residual and why it isn't lower
  • Signal injection test in the FSS EOM path to see bad conversion of phase actuation in to intensity
  • AEOM phase conversion

Need to make thermal hats for all the remaining Lithium Niobate based modulators and maybe even the wave plates.

 

Quote:

There is a big noise bump at 200-3000 Hz. Its not explained by the noise budget. How come?

Make some hypothesis an then test them.

 

ELOG V3.1.3-